# Analysis of the near-side ridge structure in pp collisions via Momentum-Kick Model

### v1

@@ III.3 Extension to Other Multiplicity Ranges
First, to set the parameter $q$, we use its inverse proportional relationship to the multiplicity, as found in Subsect III.2:
@@
Second, since the temperature reflects the statistical average of kinetic motion, we adjust it based on the $\langle p_{T}\rangle$ ratio for each multiplicity bin, using the method applied in Subsect III.2. This can be expressed using the following equation:
@@
$54.5$ 49.549.549.549.560.560.560.560.5 $1.123$ $2.237$ $1.086$ $0.221$ $0.094$
$63.58$ 59.559.559.559.570.570.570.570.5 $1.133$ $1.917$ $1.105$ $0.047$ $0.297$
$73.48$ 68.568.568.568.580.580.580.580.5 $1.165$ $1.659$ $1.125$ $0.044$ $0.562$
$83.47$ 78.578.578.578.590.590.590.590.5 $1.185$ $1.460$ $1.144$ $0.051$ $0.982$
$93.38$ 88.588.588.588.5100.5100.5100.5100.5 $1.200$ $1.305$ $1.160$ $0.064$ $1.645$
$104.8$ 100.5100.5100.5100.5110.5110.5110.5110.5 $1.216$ $1.163$ $1.178$ $0.061$ $2.020$
$114.8$ 110.5110.5110.5110.5120.5120.5120.5120.5 $1.241$ $1.062$ $1.203$ $0.082$ $3.249$
$124.7$ 120.5120.5120.5120.5130.5130.5130.5130.5 $1.259$ $0.977$ $1.220$ $0.111$ $5.112$
$137.4$ 130.5130.5130.5130.5150.5150.5150.5150.5 $1.263$ $0.887$ $1.224$ $0.111$ $6.028$
Table 4: The results of the physical parameters and the values used for each multiplicity range. The first column represents the average multiplicity of the multiplicity range shown in the second column. The sixth column represents the overall constant ‘$A$’ in $f_{R}\langle N_{k}\rangle$, and the last column represents $f_{R}\langle N_{k}\rangle$ itself. The second row shows parameters found in the MKM using an overall high-multiplicity range fitted for ALICE and CMS in Subsect III.2. $q$ and $T$ values in various multiplicity bins are calculated from Eq. (9)–(11) based on these values. The last five rows represent the high-multiplicity part defined by the ATLAS [50].
The $y$-axis ($Y_{\text{int}}$) represents the integrated near-side yield, while the $x$-axis ($N_{\text{ch}}^{\text{rec}}$) signifies the multiplicity. The black dots denote $Y_{\text{int}}^{\text{ZYAM}}$, the integrated near-side yield using the ZYAM method. We observed that the overall yield of near-side particles increases with the multiplicity.
Consequently, we introduced linear multiplicity dependence into $f_{R}\langle N_{k}\rangle$, following the blue solid line in Figure 5, which is
$f_{R}\langle N_{k}\rangle=A(-35+0.65\langle N_{\text{ch}}^{\text{rec}}\rangle).$ (11)
@@
Figure 7: Predictions of the $\Delta\phi$ correlation at 14 TeV from the MKM for CMS (black), ALICE (red), and ATLAS (blue). The ALICE and CMS predictions are plotted for $41<p_{T}<$ 4 GeV, while ATLAS is plotted for $50.5<p_{T}<$ 5 GeV. These predictions are based on the parameters extrapolated from the 7 TeV and 13 TeV results in Subsect III.2. The parameter values are adjusted by ±$\pm 10\%$ to determine upper and lower bounds.
IV Prediction
@@
At this collision energy, we do not have the multiplicity and $\langle p_{T}\rangle$ values needed to predict the key parameters of the MKM. Therefore, we apply the energy ratio between high-multiplicity pp collisions at 13 TeV and 7 TeV, as obtained from Subsection III.2, to extend our predictions to 14 TeV. This approach is based on the assumption that the multiplicity and $\langle p_{T}\rangle$ are proportional to the center-of-mass energy. We adjust parameter values by ±$\pm 10\%$ to derive the upper and lower bounds. Table 5 shows the predicted parameter values for pp collisions at 14 TeV, and Figure 7 illustrates the MKM’s prediction for the $\Delta\phi$ correlation in high-multiplicity pp collisions at 14 TeV. Three plots represent the CMS (black), ALICE (red), and ATLAS (blue) experimental conditions, respectively.
@@
V CONCLUSIONS
@@
In conclusion, the MKM has successfully described the ridge structure in small systems and is expected to be applicable to future LHC results. To predict the pp collision at 14 TeV data, we extrapolated parameter values from results at 7 TeV and 13 TeV. In order to establish the upper and lower bounds, we adjusted parameter values by ±$\pm 10\%$. This prediction result can be tested through future experiments in the LHC.

### v2

@@ 3.3 Extension to Other Multiplicity Ranges
First, to set the parameter $q$, we use its inverse proportional relationship to the multiplicity, as found in Subsect 3.2:
@@
Second, since the temperature reflects the statistical average of kinetic motion, we adjust it based on the $\langle p_{T}\rangle$ ratio for each multiplicity bin, using the method applied in Subsect 3.2. This can be expressed using the following equation:
@@
$54.5$ 49.549.549.549.560.560.560.560.5 $1.123$ $2.237$ $1.086$ $0.082$ $0.118$
$63.58$ 59.559.559.559.570.570.570.570.5 $1.133$ $1.917$ $1.105$ $0.265$ $0.306$
$73.48$ 68.568.568.568.580.580.580.580.5 $1.165$ $1.659$ $1.125$ $0.523$ $0.571$
$83.47$ 78.578.578.578.590.590.590.590.5 $1.185$ $1.460$ $1.144$ $0.926$ $0.980$
$93.38$ 88.588.588.588.5100.5100.5100.5100.5 $1.200$ $1.305$ $1.160$ $1.604$ $1.665$
$104.8$ 100.5100.5100.5100.5110.5110.5110.5110.5 $1.216$ $1.163$ $1.178$ $1.953$ $2.021$
$114.8$ 110.5110.5110.5110.5120.5120.5120.5120.5 $1.241$ $1.062$ $1.203$ $3.202$ $3.277$
$124.7$ 120.5120.5120.5120.5130.5130.5130.5130.5 $1.259$ $0.977$ $1.220$ $5.062$ $5.144$
$137.4$ 130.5130.5130.5130.5150.5150.5150.5150.5 $1.263$ $0.887$ $1.224$ $5.949$ $6.038$
Table 4: The results of the physical parameters and the values used for each multiplicity range. The first column represents the average multiplicity of the multiplicity range shown in the second column. The sixth column represents the overall constant ‘$A$’ in $f_{R}\langle N_{k}\rangle$, and the last column represents $f_{R}\langle N_{k}\rangle$ itself. The second row shows parameters found in the MKM using an overall high-multiplicity range fitted for ALICE and CMS in Subsect 3.2. $q$ and $T$ values in various multiplicity bins are calculated from Eq. (9)–(11) based on these values. The last five rows represent the high-multiplicity part defined by the ATLAS [3].
The $y$-axis ($Y_{\text{int}}$) represents the integrated near-side yield, while the $x$-axis ($N_{\text{ch}}^{\text{rec}}$) signifies the multiplicity. The black dots denote $Y_{\text{int}}^{\text{ZYAM}}$, the integrated near-side yield using the ZYAM method. We observed that the overall yield of near-side particles increases with the multiplicity. Consequently, we introduced linear multiplicity dependence into $f_{R}\langle N_{k}\rangle$, following the blue solid line in Figure 7, which is
$f_{R}\langle N_{k}\rangle=A+0.00065\langle N_{\text{ch}}^{\text{rec}}\rangle.$ (11)
@@
4 Prediction
@@
At this collision energy, we do not have the multiplicity and $\langle p_{T}\rangle$ values needed to predict the key parameters of the MKM. Therefore, we apply the energy ratio between high-multiplicity pp collisions at 13 TeV and 7 TeV, as obtained from Subsection 3.2, to extend our predictions to 14 TeV. This approach is based on the assumption that the multiplicity and $\langle p_{T}\rangle$ are proportional to the center-of-mass energy. We adjust parameter values by ±$\pm 10\%$ to derive the upper and lower bounds. Table 5 shows the predicted parameter values for pp collisions at 14 TeV, and Figure 9 illustrates the MKM’s prediction for the $\Delta\phi$ correlation in high-multiplicity pp collisions at 14 TeV. Three plots represent the CMS (black), ALICE (red), and ATLAS (blue) experimental conditions, respectively.
@@
Figure 9: Predictions of the $\Delta\phi$ correlation at 14 TeV from the MKM for CMS (black), ALICE (red), and ATLAS (blue). The ALICE and CMS predictions are plotted for $41<p_{T}<$ 4 GeV, while ATLAS is plotted for $50.5<p_{T}<$ 5 GeV. These predictions are based on the parameters extrapolated from the 7 TeV and 13 TeV results in Subsect 3.2. The parameter values are adjusted by ±$\pm 10\%$ to determine upper and lower bounds.
5 CONCLUSIONS
@@
In conclusion, the MKM is a very simplified model with few parameters through scattering but has successfully described the ridge structure in small systems. It is expected to be applicable to future LHC results. To predict the pp collision at 14 TeV data, we extrapolated parameter values from results at 7 TeV and 13 TeV. In order to establish the upper and lower bounds, we adjusted parameter values by ±$\pm 10\%$. This prediction result can be tested through future experiments in the LHC.